# A Survey on Self-play Methods in Reinforcement Learning

### v3

@@ 3.3.5 R-NaD
Although R-NaD (Perolat et al., 2021) is initially described as leveraging evolutionary game theory with a regularization component. Here, we categorized it into the PSRO series with a unique oracle computation technique (Algo. 3), executed in two stages: the first stage involves transforming the reward based on the regularization policy to make it policy-dependent, and the second stage applies replicator dynamics (Taylor & Jonker, 1978) until convergence to a fixed point. The oracle added to the policy population $\Pi$ in each iteration is derived from the reward-transformed game, not the original problem’s oracle. Nonetheless, this approach ensures that the policy converges to the NE of the original game when the game is monotone. The MSS of R-NaD is the vanilla MSS, as described in Equ. (5). This equation illustrates that the fixed point reached in each iteration, the oracle, is utilized as the regularization policy for the next iteration.
3.4 Ongoing-training-based Series
@@
3.4.1 Integration into Our Framework
We can incorporate these ongoing-training-based series of algorithms into our proposed framework (Algo. 1) using the following settings: First, we use immediate initialization to initialize $\Pi$ because, in the ongoing-training-based series, all policies in the policy population are trained together, rather than the policy population growing with each iteration. Second, we set $E=E_{max}>1$, which represents the maximum number of epochs to optimize each policy within the policy population. In other words, each unique policy undergoes iterative training for $E_{max}$ times. Third, since each policy undergoes training for $E_{max}$ times, we utilize $\pi_{i}(\cdot|h(i))$ to initialize $\pi_{i}^{h}$. This means that policy updates are self-referential. Lastly, under Assumption 1, different from Collory 1 and 2, due to the continuous training process of all policies, we derive Collory 3.
@@ 3.4.1 Integration into Our Framework
When policy $i$ is selected for training, policy $k$ $(k\geq i)$ has already been actually initialized and is therefore considered an active policy. Furthermore, in epoch $e$ $(e>1)$, policy $k$ has already been updated and holds significant meaning. As a result, policy $k$, where $k\geq i$, is likely to be chosen as the opponent for policy $i$. Consequently, the interaction matrix $\Sigma$ is generally not a lower triangular matrix. ∎
3.4.2 FTW
@@
Quake III Arena Capture the Flag is a renowned 3D multi-player first-person video game where two teams vie to capture as many flags as possible. The For The Win (FTW) agent (Jaderberg et al., 2019) is designed to perform at human-level proficiency in this game. A pivotal aspect of FTW is its employment of the ongoing-training-based self-play in RL. Specifically, it trains a population of $N$ different policies in parallel, which compete and collaborate with each other. When policy $i$ is undergoing training, FTW samples both its teammates and adversaries from the population. Specially, in scenarios where each team comprises a single member, it can be seamlessly integrated into our framework using the subsequent MSS:
$\textit{MSS}(\mathcal{P})_{ij}=\frac{1}{N}.$ (10)
@@
This essentially means that the interaction graph is densely connected. Moreover, all policies draw upon a unified policy network parameterized by $\phi$. Hence, $\pi_{i}(\cdot|h(i))$ can be aptly depicted as $\pi_{\phi}(\cdot|h(i))$. Furthermore, since these policies are not conditioned on external parameters, it is straightforward to represent the conditioning function $h(i)=\emptyset$.
3.4.3 NeuPL
@@
Neural Population Learning (NeuPL) (Liu et al., 2022c) introduces another critical innovation: it employs a unified conditional network, where each policy is adjusted against a specific meta-game mixture strategy. This is instrumental in enabling transfer learning across policies. Owing to NeuPL’s reliance on a unified conditional network parameterized by $\theta$, $\pi_{i}(\cdot|h(i))$ can be succinctly represented as $\pi_{\theta}(\cdot|h(i))$. Since NeuPL policies depend on the opponent sampling strategy $\sigma_{i}$, we define $h(i)=\sigma_{i}$.
3.4.4 Simplex-NeuPL
@@
Simplex-NeuPL (Liu et al., 2022a), which builds on NeuPL, is designed to achieve any-mixture optimality, which signifies that the formulated policy should exhibit flexibility across a diverse spectrum of adversaries, including those that might not present equivalent competitive prowess. To model the population learning process from a geometric perspective, Simplex-NeuPL introduces the concept of the population simplex. Analogously to its predecessor, Simplex-NeuPL integrates a conditional network to characterize the policy, represented as $\pi_{\theta}(\cdot|h(i))$ conditioned on the opponent sampling strategy $h(i)=\sigma_{i}$. Intriguingly, there is a slight possibility that $\sigma_{i}$ does not originate from the MSS. Instead, it is drawn as a sample from the population simplex. This sampling mechanism results in greater robustness.
3.5 Regret-minimization-based Series
@@
Another line of self-play algorithms is based on regret minimization. Unlike other categories, this series prioritizes accumulated payoff over time rather than focusing solely on a single episode, which is common in repeated games. For example, in games like Texas Hold’em or Werewolf, players must use deception, concealment, and bluffing to aim for overall victories rather than just winning a single game. Traditional regret-minimization-based self-play typically doesn’t use RL, but later studies have combined them for improved performance. This section will also discuss traditional regret-minimization methods to lay the groundwork for understanding their integration with RL.
3.5.1 Integration into Our Framework
@@
We can also integrate the regret-minimization-based series of algorithms into our proposed framework (Algo. 1) with the following settings: First, similar to traditional self-play algorithms and the PSRO series, we use lazy initialization to initialize the policy population $\Pi$. Second, we set $E=1$, and $N$ is regarded as the maximum iteration to optimize the policy. Third, we initialize $\pi_{i}^{h}$ using $\pi_{i-1}(\cdot|h(i-1))$ to utilize the most current trained policy. More specifically, $h(i)=h(i-1)$ and $\pi_{i}^{h}=\pi_{i-1}^{h}$. Fourth, in each iteration $i$, $h(i)$ represents the specific elements that regret-minimization-based self-play algorithms need to store. Note that this category relies heavily on the information in $h(i)$. For instance, in Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007), it is necessary to store counterfactual regrets for every action in every information set for each player within $h(i)$. Once $h(i)$ is determined, the corresponding policy is also defined. We will discuss CFR in detail in Sec. 3.5.2. Fifth, the ABR operator (Algo. 4) incorporates regret-related information into $h(i)$. Unlike the original CFR, Algo. 4 updates the regrets of each player sequentially. This means that the regrets of player 2 are updated after considering the already-updated regrets of player 1. This adjustment has been shown to accelerate convergence empirically and possesses a theoretical error bound (Burch et al., 2019). Additionally, each $\pi_{i}^{h}$ represents the strategies for all players, and in iteration $i$, player $j$ uses the strategy $\pi_{i}^{h}(j)$. Lastly, the MSS of the regret-minimization-based series is the vanilla MSS, as described in Equ. (5). Also, we can derive Collory 4.
@@
Regret minimization-based algorithms always use the latest strategy for training. In other words, at iteration $i$, $\pi_{i-1}^{h}$ is consistently chosen as the opponent policy. Consequently, the interaction matrix $\Sigma$ is a unit lower shift matrix. ∎
3.5.2 Vanilla CFR
@@
Regret measures the difference between the best possible payoff and the actual payoff. The regret matching algorithm (Hart & Mas-Colell, 2000) optimizes decisions by selecting strategies based on accumulated positive overall regrets, with higher overall regret strategies being more likely chosen to correct past underperformance. After each round, players update the overall regret values for each strategy. Strategies will converge to an approximate NE. However, this method primarily applies to normal-form games because computing overall regret in extensive-form games is challenging.
Zinkevich et al. (2007) proposes CFR for extensive-form games. Here, it is referred to as vanilla CFR to distinguish it from later advancements. Vanilla CFR maintains the strategy and the immediate counterfactual regrets for each information set. Theoretically, this regret is specific to each information set, and the aggregation of these immediate counterfactual regrets is an upper bound for the overall regret. Therefore, the problem can be decomposed into numerous more minor regret minimization problems.
Next, there is a more detailed description of vanilla CFR. In this survey, $\pi_{i}$ denotes the strategy at iteration $i$ for consistency, whereas the original paper uses $\sigma$. Furthermore, we use $j$ to represent player $j$, and $\pi_{i}(j)$ denotes the policy used by player $j$ at iteration $i$. The counterfactual value $v_{j}(\pi,I)$, which represents the expected value upon reaching the information set $I$ when all players, except for player $j$, adhere to the strategy $\pi$:
$v_{j}(\pi,I)=\sum_{h\in I,h^{\prime}\in Z}\text{Prob}_{-j}^{\pi}(h)\text{Prob}^{\pi}(h,h^{\prime})u_{j}(h^{\prime}).$ (11)
It is an unnormalized form of the counterfactual utility in the original paper. Based on this counterfactual value, the immediate counterfactual regret is defined as:
$R_{j,\text{imm}}^{T}(I)=\frac{1}{T}\max_{a\in A(I)}\sum_{i=1}^{T}\left(v_{j}(\pi_{i}|_{I\to a},I)-v_{j}(\pi_{i},I)\right),$ (12)
where $\pi_{i}|_{I\to a}$ denotes player $j$ choose action $a$ with probability $1$ at iteration $i$. Moreover, the positive immediate counterfactual regret is:
$R_{j,\mathrm{imm}}^{T,+}(I)=\mathrm{max}(R_{j,\mathrm{imm}}^{T}(I),0).$ (13)
The player $j$’s average overall regret $R_{j}^{T}$ is defined as:
$R_{j}^{T}=\dfrac{1}{T}\max_{\pi(j)^{*}\in\Pi(j)}\sum_{i=1}^{T}\left(u_{j}(\pi(j)^{*},\pi_{i}(-j))-u_{j}(\pi_{i})\right).$ (14)
Theoretically, average overall regret $R_{j}^{T}$ is bounded by the sum of positive immediate counterfactual regrets:
$R_{j}^{T}\leq\sum_{I\in\mathcal{I}_{j}}R_{j,\text{imm}}^{T,+}(I).$ (15)
Thus, the problem can be simplified from minimizing the overall regret to minimizing the immediate counterfactual regrets in each information set separately. For simplicity, we often drop "immediate" in discussions and refer directly to counterfactual regrets. The counterfactual regret can also be recorded as follows:
$R_{j}^{T}(I,a)=\frac{1}{T}\sum_{i=1}^{T}\left(v_{j}(\pi_{i}|_{I\to a},I)-v_{j}(\pi_{i},I)\right),$ (16)
$R_{j}^{T,+}(I,a)=\max\{R_{j}^{T}(I,a),0\}.$ (17)
The regret matching algorithm (Hart & Mas-Colell, 2000) determines the strategy at each information set:
$\left.\pi_{j}^{T+1}(I)(a)=\left\{\begin{array}[]{ll}\frac{R_{j}^{T,+}(I,a)}{\sum_{a\in A(I)}R_{j}^{T,+}(I,a)}&\text{ if denominator}>0\\ \frac{1}{|A(I)|}&\text{ otherwise}\end{array}\right.\right..$ (18)
It’s worth noting that in some studies, the normalization factor $\frac{1}{T}$ is omitted in Equ. (12), (14) and (16).
@@
Vanilla CFR has several shortcomings. Firstly, it requires traversing the entire game tree in each iteration, computationally intractable for larger game trees. Although some efforts have focused on game abstraction to reduce the size of the game tree, greater abstraction can lead to decreased performance. Second, it requires storing counterfactual regrets $R^{i}(I,a)$ for every action $a$ in every information $I$ at each iteration $i$. These values are stored in $h(i)$ within our proposed framework (Algo. 1), leading to significant storage challenges.
3.5.3 Time-saving Variants of CFR
@@
Many studies focus on enhancing the time efficiency of CFR. The first approach involves modifying the regret calculation to increase its speed. CFR+ (Tammelin, 2014) implements regret-matching+ by storing non-negative regret-like values $R_{j}^{+,i}(I,a)$, rather than $R_{j}^{i}(I,a)$. Additionally, CFR+ updates the regrets of each player sequentially and adopts a weighted average strategy. Moreover, Brown & Sandholm (2019a) introduces the concept of weighted regrets and develops Linear CFR (LCFR) and Discounted CFR (DCFR). The second approach involves adopting sampling methods. While it requires more iterations, each iteration is shorter, reducing the overall convergence time. Monte Carlo CFR (MCCFR) (Lanctot et al., 2009) divides terminal histories into blocks, with each iteration sampling from these blocks instead of traversing the entire game tree. This allows for calculating sampled counterfactual values for each player, leading to counterfactual regrets that, in expectation, match those of the Vanilla CFR. Vanilla CFR is a specific case where all histories are divided into just one block. MCCFR typically manifests in three forms: outcome-sampling MCCFR, where each block corresponds to a single history; external-sampling MCCFR, which samples opponent and chance nodes; and chance-sampling MCCFR, only focusing on chance nodes. Moreover, Johanson et al. (2012) expands on chance-sampling by categorizing based on the handling of public and private chance nodes. Some studies also focus on learning how to reduce the variance of MCCFR to speed up convergence (Schmid et al., 2019; Davis et al., 2020). In addition to these two primary approaches, other studies have identified that warm starting (Brown & Sandholm, 2016) and pruning (Brown & Sandholm, 2015; Brown et al., 2017; Brown & Sandholm, 2017) can also accelerate convergence.
3.5.4 Space-saving Variants of CFR
@@
In perfect information games, decomposition reduces problem-solving scale by solving subgames. However, in imperfect information games, defining subgames is challenging due to their intersection with information set boundaries. CFR-D (Burch et al., 2014) is a pioneering method for decomposing imperfect information games into the main component called the trunk and subgames defined by forests of trees that do not divide any information sets. In each iteration, CFR is applied within the trunk for both players, and a solver is used to determine the counterfactual BR in the subgame. The process includes updating the trunk with the counterfactual values from the subgame’s root and updating the average counterfactual values at this root, while the solution to the subgame is discarded. CFR-D minimizes storage needs by only saving $R^{i}(I_{*},a)$ for information sets in the trunk and at each subgame’s root, which is denoted by $I_{*}$, trading off storage efficiency against the time required to resolve subgames. Similar thoughts are echoed in the Continue-Resolving technique used by DeepStack (Moravčík et al., 2017) and the Safe and Nested Subgame Solving technique used by Libratus (Brown & Sandholm, 2018). We will discuss these approaches in Sec. 4.2.1, exploring their application to Texas Hold’em.
3.5.5 Estimation Variants of CFR

### v4

@@ 3.3.5. R-NaD
Although R-NaD (Perolat et al., 2021) is initially described as leveraging evolutionary game theory with a regularization component. Here, we categorized it into the PSRO series with a unique oracle computation technique (Algo. 3), executed in two stages: the first stage involves transforming the reward based on the regularization policy to make it policy-dependent, and the second stage applies replicator dynamics (Taylor and Jonker, 1978) until convergence to a fixed point. The oracle added to the policy population $\Pi$ in each iteration is derived from the reward-transformed game, not the original problem’s oracle. Nonetheless, this approach ensures that the policy converges to the NE of the original game when the game is monotone. The MSS of R-NaD is the vanilla MSS, as described in Equ. (5). This equation illustrates that the fixed point reached in each iteration, the oracle, is utilized as the regularization policy for the next iteration.
3.4. Ongoing-Training-Based Series
@@
3.4.1. Integration into Our Framework
We can incorporate these ongoing-training-based series of algorithms into our proposed framework (Algo. 1) using the following settings: First, we use immediate initialization to initialize $\Pi$ because, in the ongoing-training-based series, all policies in the policy population are trained together, rather than the policy population growing with each iteration. Second, we set $E=E_{max}>1$, which represents the maximum number of epochs to optimize each policy within the policy population. In other words, each unique policy undergoes iterative training for $E_{max}$ times. Third, since each policy undergoes training for $E_{max}$ times, we utilize $\pi_{[k]}(\cdot|h(k))$ to initialize $\pi_{[k]}^{h}$. This means that policy updates are self-referential. Lastly, under Assumption 1, different from Collory 3.1 and 3.2, due to the continuous training process of all policies, we derive Collory 3.3.
@@ 3.4.1. Integration into Our Framework
When policy $m$ is selected for training, policy $n$ $(n\geq m)$ has already been actually initialized and is therefore considered an active policy. Furthermore, in epoch $e$ $(e>1)$, policy $n$ has already been updated and holds significant meaning. As a result, policy $n$, where $n\geq m$, is likely to be chosen as the opponent for policy $m$. Consequently, the interaction matrix $\Sigma$ is generally not a lower triangular matrix. ∎
3.4.2. FTW
@@
Quake III Arena Capture the Flag is a renowned 3D multi-player first-person video game where two teams vie to capture as many flags as possible. The For The Win (FTW) agent (Jaderberg et al., 2019) is designed to perform at human-level proficiency in this game. A pivotal aspect of FTW is its employment of the ongoing-training-based self-play in RL. Specifically, it trains a population of $K$ different policies in parallel, which compete and collaborate with each other. When policy $k$ is undergoing training, FTW samples both its teammates and adversaries from the population. Specially, in scenarios where each team comprises a single member, it can be seamlessly integrated into our framework using the subsequent MSS:
(10) $\textit{MSS}(\mathcal{P})_{mn}=\frac{1}{K}.$
@@
This essentially means that the interaction graph is densely connected. Moreover, all policies draw upon a unified policy network parameterized by $\phi$. Hence, $\pi_{[k]}(\cdot|h(k))$ can be aptly depicted as $\pi_{\phi}(\cdot|h(k))$. Furthermore, since these policies are not conditioned on external parameters, it is straightforward to represent the conditioning function $h(k)=\emptyset$.
3.4.3. NeuPL
@@
Neural Population Learning (NeuPL) (Liu et al., 2022c) introduces another critical innovation: it employs a unified conditional network, where each policy is adjusted against a specific meta-game mixture strategy. This is instrumental in enabling transfer learning across policies. Owing to NeuPL’s reliance on a unified conditional network parameterized by $\theta$, $\pi_{[k]}(\cdot|h(k))$ can be succinctly represented as $\pi_{\theta}(\cdot|h(k))$. Since NeuPL policies depend on the opponent sampling strategy $\sigma_{[k]}$, we define $h(k)=\sigma_{[k]}$.
3.4.4. Simplex-NeuPL
@@
Simplex-NeuPL (Liu et al., 2022a), which builds on NeuPL, is designed to achieve any-mixture optimality, which signifies that the formulated policy should exhibit flexibility across a diverse spectrum of adversaries, including those that might not present equivalent competitive prowess. To model the population learning process from a geometric perspective, Simplex-NeuPL introduces the concept of the population simplex. Analogously to its predecessor, Simplex-NeuPL integrates a conditional network to characterize the policy, represented as $\pi_{\theta}(\cdot|h(k))$ conditioned on the opponent sampling strategy $h(k)=\sigma_{[k]}$. Notably, $\sigma_{[k]}$ may be sampled from the population simplex rather than the MSS, enhancing robustness.
3.5. Regret-Minimization-Based Series
@@
Another line of self-play algorithms is based on regret minimization (RM). Unlike other categories, this series aims to minimize cumulative regret over time, a concept rooted in online learning theory. This distinction is crucial in two-player zero-sum games, where these methods enjoy strong theoretical guarantees. Traditional regret-minimization-based self-play typically doesn’t use RL, but later studies have combined them for improved performance. This section will also discuss traditional regret-minimization methods to lay the ground for understanding their integration with RL.
3.5.1. Integration into Our Framework
@@
We can also integrate the regret-minimization-based series of algorithms into our proposed framework (Algo. 1) with the following settings: First, similar to traditional self-play algorithms and the PSRO series, we use lazy initialization to initialize the policy population $\Pi$. Second, we set $E=1$, and $K$ is regarded as the maximum iteration to optimize the policy. Third, in each iteration $k$, $h(k)$ represents the specific elements that regret-minimization-based self-play algorithms need to store. Note that this category relies heavily on the information in $h(k)$. For instance, in Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007), it is necessary to store counterfactual regrets for every action in every information set for each player within $h(k)$. Once $h(k)$ is determined, the corresponding policy is also defined. We will discuss CFR in detail in Sec. 3.5.2. Fourth, we initialize $\pi_{[k]}^{h}$ using $\pi_{[k-1]}(\cdot|h(k-1))$ to utilize the most current trained policy. More specifically, $h(k)$ is initialized by h(k1)h(k-1) and $\pi_{[k]}^{h}$ is initialized by $\pi_{[k-1]}^{h}$. In each new iteration, regret-minimization information is continuously added to $h(k)$ (Line 4 in Algo. 4). As a result, $h$ accumulates information across all iterations until the end of training. Fifth, the ABR operator (Algo. 4) incorporates regret-related information into $h(k)$. Unlike the original CFR, Algo. 4 updates the regrets of each player sequentially. This means that the regrets of player 2 are updated after considering the already-updated regrets of player 1. This adjustment has been shown to accelerate convergence empirically and possesses a theoretical error bound (Burch et al., 2019). Each $\pi_{[k]}^{h}$ represents the strategies for all players, and in iteration $k$, player $i$ uses the strategy $\pi_{[k]}^{h}(i)$. Lastly, the MSS of the regret-minimization-based series is the vanilla MSS, as described in Equ. (5). Also, we can derive Collory 3.4.
@@
Regret-minimization-based algorithms always train against the latest strategy; that is, at iteration $k$, the opponent policy is $\pi_{[k-1]}^{h}$. As a result, the interaction matrix $\Sigma$ becomes a unit lower shift matrix. ∎
3.5.2. Vanilla CFR
@@
Regret measures the difference between the best possible payoff and the actual payoff. The regret matching algorithm (Hart and Mas-Colell, 2000) optimizes decisions by selecting strategies based on accumulated positive overall regrets, with higher overall regret strategies being more likely chosen to correct past underperformance. After each round, players update the overall regret values for each strategy. The algorithm is guaranteed to converge to the set of NE, and any finite stopping point yields an approximate NE. However, this method primarily applies to normal-form games because computing overall regret in extensive-form games is challenging.
Zinkevich et al. (2007) proposes CFR for extensive-form games. In this survey, we refer to it as vanilla CFR to distinguish it from later advancements. Vanilla CFR maintains both the strategy and the counterfactual regret values for each information set. Theoretically, the sum of these local regrets provides an upper bound on the overall regret. Therefore, the problem can be decomposed into many smaller regret minimization subproblems. CFR focuses on minimizing regret locally at each information set, making the approach computationally tractable. At each iteration, the strategy is updated using regret matching, and the final strategy is obtained through strategy averaging, where the strategies played across iterations are averaged over time. This averaged strategy is proven to converge to an NE in two-player zero-sum games.
@@
Vanilla CFR has several shortcomings. Firstly, it requires traversing the entire game tree in each iteration, computationally intractable for larger game trees. Although some efforts have focused on game abstraction to reduce the size of the game tree, greater abstraction can lead to decreased performance. Second, it requires storing counterfactual regrets $R^{k}(I,a)$ for every action $a$ in every information $I$ at each iteration $k$. These values are stored in $h(k)$ within our proposed framework (Algo. 1), leading to significant storage challenges.
3.5.3. Time-Saving Variants of CFR
@@
Many studies focus on enhancing the time efficiency of CFR. The first approach involves modifying the regret calculation to increase its speed. CFR+ (Tammelin, 2014) implements regret-matching+ by storing non-negative regret-like values $R^{+,k}(I,a)$, rather than $R^{k}(I,a)$ in $k^{th}$ iteration. Additionally, CFR+ updates the regrets of each player sequentially and adopts a weighted average strategy. Moreover, Brown and Sandholm (2019a) introduces the concept of weighted regrets and develops Linear CFR (LCFR) and Discounted CFR (DCFR). The second approach involves adopting sampling methods. While it requires more iterations, each iteration is shorter, reducing the overall convergence time. Monte Carlo CFR (MCCFR) (Lanctot et al., 2009) divides terminal histories into blocks, with each iteration sampling from these blocks instead of traversing the entire game tree. This allows for calculating sampled counterfactual values for each player, leading to counterfactual regrets that, in expectation, match those of the Vanilla CFR. Vanilla CFR is a specific case where all histories are divided into just one block. MCCFR typically manifests in three forms: outcome-sampling MCCFR, where each block corresponds to a single history; external-sampling MCCFR, which samples opponent and chance nodes; and chance-sampling MCCFR, only focusing on chance nodes. Moreover, Johanson et al. (2012b) expands on chance-sampling by categorizing based on the handling of public and private chance nodes. Some studies also focus on learning how to reduce the variance of MCCFR to speed up convergence (Schmid et al., 2019; Davis et al., 2020). In addition to these two primary approaches, other studies have shown that warm starting (Brown and Sandholm, 2016) and pruning (Brown and Sandholm, 2015; Brown et al., 2017; Brown and Sandholm, 2017) can also accelerate convergence.
3.5.4. Space-Saving Variants of CFR
@@
In perfect information games, decomposition reduces problem-solving scale by solving subgames. However, in imperfect information games, defining subgames is challenging due to their intersection with information set boundaries. CFR-D (Burch et al., 2014) is a pioneering method for decomposing imperfect information games into the main component called the trunk and subgames defined by forests of trees that do not divide any information sets. In each iteration, CFR is applied within the trunk for both players, and a solver is used to determine the counterfactual BR in the subgame. The process includes updating the trunk with the counterfactual values from the subgame’s root and updating the average counterfactual values at this root, while the solution to the subgame is discarded. CFR-D minimizes storage needs by only saving $R^{k}(I_{*},a)$ for information sets in the trunk and at each subgame’s root, which is denoted by $I_{*}$, trading off storage efficiency against the time required to resolve subgames. Similar thoughts are echoed in the Continue-Resolving technique used by DeepStack (Moravčík et al., 2017) and the Safe and Nested Subgame Solving technique used by Libratus (Brown and Sandholm, 2018). We will discuss these approaches in Sec. 4.2.1, exploring their application to Texas Hold’em.
3.5.5. Estimation Variants of CFR